OpenStudy (anonymous):

Multiply and simplify by factoring fourth root 16 * fourth root 128

OpenStudy (anonymous):

\[16=2^4\] so \[\sqrt[4]{16}=2\] is a start

OpenStudy (anonymous):

yea I got that 2 lol but the second part threw me off

OpenStudy (anonymous):

\[128=2^7\] so \[\sqrt[4]{128}=2\sqrt[4]{2^3}\]

OpenStudy (anonymous):

this is how i did it in my head. the exponent is 7, the "index" is 4 4 goes in to 7 one time so a 2 comes out the remainder is 3, so a 2^3 stays in

OpenStudy (anonymous):

yea that makes sense

OpenStudy (anonymous):

that explains this line \[\sqrt[4]{128}=\sqrt[4]{2^7}=2\sqrt[4]{2^3}=2\sqrt[4]{8}\]

OpenStudy (anonymous):

so your 'final answer' is \[2\times 2\sqrt[4]{8}=4\sqrt[4]{8}\]

OpenStudy (anonymous):